UT Sep 13 2022 - Night #11 of Restart - Split night w/PEPSI-PETS




Night Info

  • Sunset 18:30
  • Nautical Twilight Ends 19:23
  • Astro Twilight Ends 19:53
  • Handover to PEPSI-PETS 20:50
  • Handover from PEPSI-PETS to DEC ~01:21
  • Astro Twilight Begins 04:39
  • Nautical Twilight Begins 05:08
  • Sunrise 06:02
  • Moonrise 20:02
  • Moon illumination ~92%

Staff

  • OSA: J.Williams
  • ISp: C. Bolyard
  • Mtn Manager: P.Hartley
  • ISA/SO: A.Cardwell, B.Rothberg, R.T.Gatto
  • Technical Expert: J.Hill
  • AO:

Night Plan

ALTA prediction for Tonight

20220912_see_evol_time_Height_20_20000.png

Starting Configuration LUCIs

LUCI Checkout

  • Collect transform data - (2hrs) - confirmatory - Requires LUCIs - transforms done and will be installed before sunset
    • update oaccontrol.conf with new results from D.Miller
      • oac@oac: /lbt/oacontrol/current/etc/oacontrol.conf
      • backup file first
      • add the snippets from D.Miller in the correct place for AGW1 and AGW2
      • need to restart PCS and GCS to take hold
    • 1.2" or better seeing needed
    • Target Fields - select one best visible field
      • 21 37 31 +00 31 55 Stone_O_m12.list
      • 22 35 49 +01 31 33 Stone_P1.list
      • 00 54 02 -00 32 34 Stone_A3.list
    • Note the LUCI frame numbers associated with each offset in IDL script
    • Can be done wth LUCI SW in crippled state (does not need flex comp, remember to load EmptyQueue.xml)
  • If Transform cannot be done (conditions too poor)
    • LUCI Scripts:
      • LUCI_2022_Restart_V1.80.xml - all scripts (except LS Rot ZP)
      • Rot_ZP_V2 script
      • LUCI_Phot_ZP_OT_V6.xm - LUCI photometric standards (fixes AGW patrol problem)
    • Imaging check (1hr) - medium priority - if short on-time skip
      • ANY seeing OK
      • dither in-field
        • NGC 7027 - cycles through all filters (4pt dither) - visible until 4am - only need to run 4 dithers
      • dither out-of-field (run one)
        • M52 - JHK, dither 5' out of field - visible all night
        • M40 - Broad filters, few narrow filters - visible all night

Telescope Wakeup Continued

  • After Reconfiguration from PEPSI-PETS
  • Conifrmatory Field Aberration (2 hrs) - Critical (1hour) - Top Priority - update installed already
    • Analysis Completed
    • RFBG
      From 45 Astigmatism Measurement
      On-Sky Arcseconds
      -----------------------------

      --------------
      Tip = 4.3", Delta IE = -4.3"
      Tilt = 24.9", Delta CA = -24.9"

      rfbg
      From 0 Astigmatism Measurement
      On-Sky Arcseconds
      -----------------------------

      --------------
      Tip = -3.3", Delta IE = 3.3"
      Tilt = 27.3", Delta CA = -27.3"

      I suggest Tip = 0, Tilt = 25

      LFBG

      => Field Data Results
      Measured bi-nodal astigmatism

      lfbg
      From 45 Astigmatism Measurement
      On-Sky Arcseconds
      -----------------------------

      --------------
      Tip = 19.4", Delta IE = -19.4"
      Tilt = 16.9", Delta CA = -16.9"


      lfbg
      From 0 Astigmatism Measurement
      On-Sky Arcseconds
      -------------------------------------------
      Tip = 7.4", Delta IE = -7.4"
      Tilt = 18.0", Delta CA = -18.0”

      I suggest Tip =15, Tilt = 17 (or non at all…might be in the noise)
      • add the values to the global offsets in the M2 collimation files: /lbt/tcs/current/psf/configuration/PSF
      • SXSMLUCIFERColllimation.dat, then propagate to all the others with .csh script.
    • Can be done in parallel (data collection, analysis in serial)
    • need both AGW1 and AGW2 on the telescope and checked to execute - DOES NOT NEED LUCIs - Just AGW 1 and AGW 2
    • 1.2" or better seeing (1" or better ideal)
    • Target Fields:
      • BS9177 18:39:33.7 +49:05:38.0
      • StoneP1 22:37:45.0 -01:13:40.45
      • BS9188 23:30:33.4 +38:18:57.0
      • BS9107 03:32:03.0 +37:20:40.0

LBC Chcekout

  • Reconfigure to LBCs by 03:21 local time, unless Field Aberration or AGW Transforms are still TBD
  • TMS Testing - 1 hours - High Priority - need to make sure TMS operates on-sky OK
    • Any field in OT that has plenty of stars
  • Superfocus sequence - only if time
    • Requires 1.2" or better stable seeing
  • Photometric Standards - backup - OR - use with TMS testing if can
    • any seeing, must be clear
  • Skyflats if clear


Night Log

Googledoc link

Closed Dome

  • UT 00:49 - M1 mirror issues - currently being checked
  • UT 00:51 - bringing up LUCI GUIs on OBS2 - after cleaning out a bunch of defunct sessions
    • Iinit all ok
    • putting LUCI-2 LS back into cabinet, then will put N30 Fieldstop into FPU
    • taking L1 test exposures - 2.51 sec blind
    • L1 0001-0002
    • putting N375 camera into LUCI-1
    • taking L2 test exposures
    • L1 0001-0002
    • putting L1 N30fieldstop in turnout to check fieldstop N375 (ignore L1 0003-0004)
    • looks OK
    • checking L2 - slightly off (L2 0003-0007) - about as good as you can get
    • OBS-2 back to being sluggish again
  • UT 01:15 - LUCIs ready
  • SX M1 - several actuators problematic
  • UT 01:46 - Closed for clouds and humidity
  • UT 03:45 - switching to PEPSI-PFU
  • UT 03:50 - G.Eychaner fixed (mostly) launching LUCI GUIs on OBS-32bit, test exposures taken. However, can't launch the RTD from the GUI. - can launch manually
  • UT 08:15 - reconfiguring for bino-LBCs. IT 8716 - AGW4 WFS images giving an error, I.Ilyin cannot take test exposures - updating IT.
  • UT 08:17 - continuing with LUCI Flexcomp data collection
  • UT 11:01 - DONE FOR THE NIGHT

Closed Dome Work

  • L1_N180_MIRR L2_N180_MIRR FlexComp _Calib15.xml
  • DONE L1_N375_MIRR L2_N375_MIRR FlexComp _Verify27d.xml use PFC ON, @EL=79,57,35
  • Scripts Located: /home/lbto/dthompson/LUCI/Flexure_Compensation
  • Contact D.Thompson if any questions.

-- BarryRothberg - 11 Sep 2022
Topic revision: r9 - 06 Dec 2022, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback